OpenStudy (anonymous):

If F(x1) + f(x2) = f(x1+x2) for all real numbers x1 and x2, which of the following could define f? a) F(X) = x+1 b) f(x) = 2x c) f(x) = 1/x d) f(x) = e^X e) f(X) = x^2

OpenStudy (amistre64):

thats part of the definition of a linear function

OpenStudy (amistre64):

of course, lower math classes like to muddy affine functions with linears. an affine function is a linear, plus a nonzero constant

OpenStudy (anonymous):

Hi, sorry could you go into more detail, I'm still not catching on. I don't think i've ever heard of affine functions?

OpenStudy (amistre64):

have you heard of a linear function?

OpenStudy (amistre64):

and if you look at the options, youll notice that one of them has a constant add to it .... which might correlate with something ive posted

OpenStudy (amistre64):

otherwise, you can try out the rule that they defined for it ... f(x+y) = f(x) + f(y)

OpenStudy (anonymous):

D right? E is the constant?

OpenStudy (amistre64):

um, no. nothing is being added in that option. and "e" is a number like "pi"

OpenStudy (amistre64):

\[e^{x+y}=e^x~e^y\]therefore that doent follow the rule defined

OpenStudy (anonymous):

Oh I get it now 2 is the constant, that means that B is the answer right?

OpenStudy (amistre64):

B is the answer :) B is the function of a line that passes thru the origin (its a linear function) the affine function is A, its a linear function (+1)

OpenStudy (anonymous):

Thanks so much!! :)